# “There Has To Be a Lot That We’re Missing”: Moderating AI-Generated Content on Reddit

### v4

@@ 5.1. Community Values Shape Attitudes Towards AIGC
Our study reveals a spectrum of attitudes towards AIGC that seems to vary with the purpose and size of communities. Our exploratory crawl of subreddit rules (3.1) found that 4% of all listed subreddits had some rule governing the use of AI, but among the largest 1% of subreddits this number increased to 20%. This trend is consistent with earlier empirical work that has studied rules on Reddit and found community popularity to be the primary predictor of whether or not a subreddit has rules (FieslerEtal2018rrc). Our conversations with moderators suggest that certain types of communities, such as empathic communities (preece_empathic_1999) and  knowledge communities (jenkins_introduction_2006) may feel the most ideologically opposed to AIGC. We also see some examples of  communities of practice (lave1991situated) objecting to AIGC, but it is also likely that new communities of practice will emerge specifically devoted to AI-reliant practices [^8] https://www.reddit.com/r/aiArt/. On the whole, our findings align with those of recent work on the impact of generative AI on online communities (burtch_consequences_2024; wei2024understandingimpactaigenerated), which suggests that the social value of online communities may make them more important than ever in the age of generative AI. Platform designers and community stewards should focus on ways to increase this social value.
It will be interesting to observe how community norms towards AIGC evolve over time. Our interviews and analysis started less than a year after the launch of ChatGPT. Our findings indicate that community norms about AIGC have been in flux since this launch and will likely continue to change. It is possible that as people become more familiar with generative AI they will come to see it as compatible with their values, no different from other automated composition tools like auto-complete or spellcheck, which is an argument that we mentioned hearing in 4.1.1. If communities are to remain relevant, moderators will need to stay abreast of their community members’ stance on this evolving technology; ultimately, if community member and moderator perspectives on AIGC use are not aligned, Reddit’s distributed moderation system allows users to “agree to disagree” by leaving and finding other communities that share their stance (grimmelmann2015virtues).
At the time of writing, some platforms have begun incorporating generative AI into their products (lunden_linkedin_2023), likely in an effort to increase engagement. This may be aligned with platforms’ profit motives, but our interviews suggest that this is likely not aligned with the will of all platform users. As we mention in 4.1.3, generative AI may bring certain benefits to online communities, but given the strong current objections that we’ve surfaced, any attempts by platforms to integrate generative AI should be done via a collaborative design process with the affected communities (Gilbert20; Matias16; Matias18).
@@ 5.2. The Challenges of Establishing and Enforcing Norms
Our findings relate to and reflect on research about the power of norms to shape behavior in online communities (chancellor18; Chandrasekharan18; kiesler11; Lessig99) Platform designers may be able to aid community moderators with the work of establishing and enforcing community norms regarding AIGC.
@@ 5.2.1. Clarifying Norms in a Moment of Flux
Our interviews show that self-governing online communities are already adapting to the phenomenon of AIGC by clarifying their norms around its use and codifying these into rules. Prior research has shown that it is common for online communities to respond to sudden changes by making rules (SeeringEtal2019mec) and that this can be an effective strategy for navigating challenging transitions (Kiene16). Moderators sometimes struggle to identify AIGC, which makes enforcing rules about it a challenge; even so, discussing and clarifying rules can help inform the behavior of well-intentioned community members. It is likely that as generative AI grows in popularity and sophistication communities will continue to need to adapt their positions and practices in response. Our findings give us hope that communities will be able to successfully adapt, though it will require substantial effort on the part of both moderators and community members.
Given that AIGC is an emergent phenomenon, community norms around its use are still in flux. Community members and admins may have different notions about whether its use is acceptable or not, and this misalignment (Koshy23) may result in legitimate users inadvertently running afoul of their communities. Online communities may be able to lessen the threat of misalignment in the community by making their norms more explicit and known to their members, which has been shown to be effective with other moderation challenges (matias2019preventing). This approach is supported by our participants’ reports that the amount of AIGC has decreased since they enacted their rules banning it, although these reports are based on the moderators’ perceptions, and there are no reliable ways to independently verify them.
@@
The importance of communicating norms about AIGC use also has implications for platform designers who may devise additional ways to ensure well-meaning posters are aware of community rules. For example, a community’s rules could be displayed in the authoring UI, perhaps with a checkbox that users must mark to indicate that they have read and are abiding by the rules. This is similar to how privacy policies are commonly presented. However, prior work has shown that privacy policies are rarely read (obar_biggest_2020), which suggests that adding a checkbox to the posting experience might only have the effect of adding friction to the experience and decreasing post volume. Another option would be to display the rules as placeholder text in the UI element where posts are written, which a previous study suggested as an elegant way to remind users of norms at the time of posting (Kiene16). That being said, displaying rules in a creative way may not be enough, as past research has demonstrated that many posters to online communities simply do not read the community’s rules before posting (JhaverAppling19). Communities concerned about this could also consider more thorough “onboarding experiences” to ensure new members are familiar with community rules, such as an interview with an existing member. However this approach places an additional time burden on both the existing community and the new member and may deter potential members who are wary of such a commitment.
Some social platforms are currently experimenting with incorporating generative AI features into their authoring tools (lunden_linkedin_2023). We suspect that the low effort required to use embedded generative AI tools will increase the overall amount of AIGC on platforms that add them. In order to respect the autonomy of the communities that sustain them, platforms should give individual communities the ability to decide for themselves whether or not to enable platform-provided generative AI tools. If a community opts out, this decision could be communicated to posters via the UI, perhaps by displaying the tools in a disabled state with an informative tooltip. Communicating the community’s stance on AIGC in this way may have the added benefit of also discouraging posting AIGC produced by off-platform generative AI tools.
@@ 5.2.2. Equitably Enforcing AIGC Bans
Our findings detail some of the challenges that moderators currently experience enforcing AIGC restrictions and bans. These challenges impact both moderators and community members, and our findings may help offer potential solutions.
@@
Even before having to enforce rules about AIGC, moderators were stretched thin trying to protect their communities from problematic content (e.g. mis- and disinformation) and behavior (e.g. harassment, incivility, and influence campaigns). With detection of AIGC being an unsolved technical problem (edwards23), moderators are left to design and use heuristics to enforce restrictions or bans. Our findings show that these heuristics can be time-intensive. Worse, the heuristics might also be flawed, or lead to biases as people’s intuitions about AIGC are often wrong in predictable ways (Jakesch_2023). Such consistent biases could disproportionately affect minorities and individuals seeking legitimate participation (liang23), such as non-native speakers, newcomers, and gender or racial minorities. It is concerning that most of our participants were confident in the efficacy of their heuristics and unaware of any potential bias. Our participants’ current reliance on potentially biased heuristics suggests the importance of both educating moderators about these biases and conducting further studies to evaluate the efficacy of detection heuristics in the community moderation context.
Platform designers could aid the work of moderators by building moderation tools that surface the relevant signals that moderators use for their detection heuristics. Earlier work has suggested tools that surface “synthesized social signals,” or “social signals computationally derived from an account’s history,” to help online community members evaluate the credibility of other accounts (Im20). Such a tool could be useful here if it additionally surfaced information about users’ interactions with the system (e.g. time spent editing a post) that would give moderators a sense of the effort that authors put into their posts, though this may be difficult to implement in a way that respects users’ privacy. Additional information about a poster’s past behavior and other community memberships may also be relevant, similar to how these are useful in assessing if a user is likely to post misinformation (Bozarth2023; diakopoulos_diamonds_2010).
Given the combination of concerns about AIGC and the lack of tools to detect it, we expect to see communities developing their own, non-technical, communication norms for demonstrating authenticity, similar to past developments such as timestamping and triforcing (bernstein14), practices used in anonymous and pseudonymous online communities such as Reddit. These techniques are a way to signal in-group membership by presenting something that only an authentic poster would be able to present. Platforms may also want to update their authoring tools to support any new community-specific signaling practices that users develop to signal their authenticity.
@@ 6. Limitations and Future Work
Our qualitative study is based on interviews with moderators of some of the largest subreddit communities, but, naturally, may not cover sentiments that generalize across the entire Reddit platform. Future work could perform a platform-wide quantitative study that could quantify trends and patterns across subreddits of different size and type. Further, since AIGC is an internet-scale phenomenon, future studies should explore how moderators are responding on other platforms besides Reddit. We recruited from popular subreddits in an attempt to target those that are currently most affected, but our participants who also moderate smaller subreddits shared that their experiences and concerns are relevant to communities of various sizes. While we took steps to verify the accurate representation of our participants’ views, several emphasized that they can only speak to their own experience and do not claim to speak on behalf of their entire communities. Because our recruitment strategy targeted subreddits with explicit rules about AIGC, we also likely omitted those communities that have fully embraced AIGC and thus have not created any rules against its use. Additional studies could include moderators from such communities for a more complete picture of the range of ways that online communities are responding to AIGC, and to further explore the relationship between a community’s values and its stance on AIGC.
Our participants mentioned heuristics that they use to detect AIGC, but we have not empirically verified their accuracy. Existing studies have demonstrated the inaccuracy of detection heuristics in a lab setting (Jakesch_2023), but further research could measure the accuracy and bias of the detection heuristics of “experts” who are working in a context that they are deeply familiar with, such as community content moderators.
@@ 7. Conclusion
We performed a qualitative investigation into volunteer community moderators on the social sharing site Reddit’s experiences with, and attitudes towards, AI-generated content use in their online communities. Rather than try to represent the range of moderators’ attitudes and experiences across all of Reddit, we focus our study on the moderators of communities who have decided to restrict the use of AIGC, in the hopes of understanding how the CSCW research community may support these efforts towards community self-determination. We find both ideological and practical objections to AIGC that make our participants feel AIGC does not belong in their communities, despite also hearing about its potential to make it easier for some community members to contribute. On the whole our participants reported concern that allowing AIGC would decrease the quality of conversation and interactions in their communities, but they felt that the current restrictions that they put in place were effective, for now, at protecting their communities. These restrictions primarily served to clarify and communicate the community norms around an emerging issue with generally unsettled norms, but the actual enforcement of these restrictions is imperfect.
As generative AI usage continues to grow, the CSCW community needs to continue to pay attention to how it affects the dynamics, practices, and health of online communities. There is much that the research community can do to develop tools and practices that may help community moderators guide their communities through this period of technological change. As a starting point, this paper explores the perspectives of moderators at this critical juncture, where, as one moderator commented, “there has to be a lot that we’re missing.”

### v5

@@ 5.1. Community Values Shape Attitudes Towards AIGC
Our study reveals a spectrum of moderator attitudes that vary with the purpose and size of communities. We expect certain objections to AIGC will lessen over time, while others will be more persistent.
5.1.1. Community Types and AIGC Attitudes
Our participants’ concerns about AIGC were largely based on the types of communities they moderated. Accordingly, it is useful to consider our results through the lens of several online community types from the Social Computing literature, such as  knowledge communities (jenkins_introduction_2006), empathic communities (preece_empathic_1999),  communities of practice (lave1991situated), and the five Community Archetypes presented by Prinster et al. (Prinster24): Topical Q&A, Learning & Perspective Broadening, Social Support, Content Generation, and Affiliation with an Entity. We suspect that AIGC’s threats to the social value of a community (Section 4.1.2) would be the most concerning to communities that emphasize authenticity (Smith23), like empathic communities (preece_empathic_1999), or communities that fit the Social Support archetype. Moderators of  knowledge communities (jenkins_introduction_2006) like r/AskHistorians and r/explainlikeimfive, which also fit the Topical Q&A archetype, also shared this concern, though they additionally took issue with AIGC’s inaccuracy (Section 4.1.1). Arguably the clearest objections came from moderators of communities of practice (lave1991situated) like r/itookapicture and r/WritingPrompts, which fit the Content Generation archetype. These communities shared both content quality and social dynamics concerns about AIGC, though we expect new communities of practice will emerge specifically devoted to AI use [^8] https://www.reddit.com/r/aiArt/. Moderators of larger, more impersonal communities that fit the Learning & Perspective Broadening archetype, such as r/news, r/CryptoCurrency, and r/politics, primarily voiced governance concerns (Section  4.1.3). We suspect this is because the size of these communities make them a target of bad actors. Finally, we primarily heard ideological objections from the communities in our sample that fit the Affiliation with an Entity archetype. We suspect this is because in our sample these are Content Generation communities (r/WoW, r/Zelda), and it remains to be seen if communities affiliated with non-media entities, like geographical places or organizations, would share similar concerns.
On the whole, our participants rejected AIGC because they felt the main purpose of their communities was to share posts created by humans—this was true across all of the community types that we encountered (of course, remember that our sampling strategy focused on communities with concerns, not representativeness). Future work should further explore the relationship between community type and attitude towards AIGC. For example, it would be interesting to see, quantitatively, whether different types of communities tend to create different types of rules about AIGC.
5.1.2. Attitudes May Shift Over Time
Our findings indicate that community norms about AIGC are still in flux. Perhaps over time people will come to treat generative AI the same as tools like auto-complete or spellcheck (a possibility we discussed in Section 4). If so, process-based objections may fade away. Of course, the opposite could happen as well: as generative AI becomes more pervasive people may seek out communities of collective refusal. In order to support community self-determination, we should design platforms that allow for both possibilities.
Additionally, generative AI will continue to change, and in time the content-based objections that we heard may no longer be relevant. For example, we consider objections based on AIGC’s accuracy and prose quality, as well as those based on unethical model training processes, to be potentially addressable by technological advances. If online communities are to remain relevant, moderators will need to stay abreast of their members’ stance on this evolving technology; ultimately, if moderator and member perspectives on AIGC differ too much, Reddit’s distributed moderation system allows users to “agree to disagree” by leaving and finding other, more aligned communities (grimmelmann2015virtues). At the time of writing, some platforms have begun incorporating generative AI into their products (lunden_linkedin_2023), likely in an effort to increase engagement. Though this might be aligned with platforms’ profit motives, our interviews suggest it is not aligned with the preferences of all users or communities, as the low effort required to use these features will likely increase the amount of AIGC on the platforms. As we mention in Section 4.1.4, generative AI might bring certain benefits, but given the strong objections that we’ve seen, attempts by platforms to integrate generative AI should be done via a collaborative design process involving users (Gilbert20; Matias16; Matias18). We suggest that platforms that host a variety of community types, such as Reddit, give individual communities the choice to opt-out of generative AI features. If communities choose to opt out, this decision could be communicated to posters via the UI, perhaps by displaying the tools in a disabled state with an informative tooltip. Signaling a community’s stance on AIGC in this way can have the added benefit of additionally discouraging posting AIGC produced off-platform. Additionally, for platforms designed for just one of the community types that we discussed, we recommend considering this community type’s specific values when deciding whether to add generative AI features. Otherwise, platforms run the risk of alienating users (Smith23).
@@ 5.2. The Challenges of Establishing and Enforcing Norms
Our findings relate to and reflect on research about the power of norms to shape behavior in online communities (chancellor18; Chandrasekharan18; kiesler11; Lessig99). We explore the specific challenges of establishing and enforcing community norms about AIGC and discuss ways in which platform designers might be able to aid community moderators with these tasks.
@@ 5.2.1. Clarifying Norms in a Moment of Flux
Our interviews show that self-governing online communities are adapting to AIGC by clarifying norms and codifying these into rules. Prior research has shown that online communities often make rules in response to sudden changes (SeeringEtal2019mec) and that this response can be an effective way to navigate challenging transitions (Kiene16). Moderators sometimes struggle to identify AIGC, which makes enforcing rules about it a challenge; even so, discussing and clarifying rules can help inform the behavior of well-intentioned community members.
@@
Community norms around AIGC are in flux and likely will be for some time. Community members and moderators may have different understandings of whether or not AIGC is acceptable, and this misalignment (Koshy23) can result in well-intentioned users inadvertently violating community norms. Communities can lessen the chance of misalignment by making norms more explicit and known to their members, which has been shown to be effective with other moderation challenges (matias2019preventing). This approach is supported by participants’ reports that AIGC has decreased since they enacted rules banning it, although these reports are based on the moderators’ perceptions, and have not been independently verified. Platform designers can help by updating UIs to ensure that well-meaning posters are aware of community rules. For example, platforms could update their authoring UIs to display a community’s rules, perhaps with a checkbox that users must mark to indicate that they have read and will abide by the rules, similar to how privacy policies are commonly presented. Another option would be to display the rules as placeholder text in the UI element where posts are written, which may remind users of norms at the time of posting (Kiene16). That said, displaying rules in a creative way may not be enough, as past research has demonstrated that privacy policies are rarely read (obar_biggest_2020) and many posters simply do not read a community’s rules before posting (JhaverAppling19). Communities may also consider more thorough “onboarding processes” to ensure new members are familiar with community rules, such as an interview with an existing member. However, this approach adds an additional time burden and might deter potential members who are wary of such a commitment.
@@ 5.2.2. Equitably Enforcing AIGC Bans
Our findings detail some of the challenges that moderators currently experience enforcing AIGC restrictions and bans. Even before the AIGC challenge, moderators were stretched thin trying to protect their communities from problematic content (e.g. mis- and disinformation) and behavior (e.g. harassment, incivility, and influence campaigns). With detection of AIGC being an unsolved problem (edwards23), moderators rely on heuristics to enforce restrictions or bans. Our findings show that these heuristics can be time-intensive. Even worse, they may be biased, as other work has shown that people’s intuitions about AIGC are wrong in predictable ways (Jakesch_2023) and that certain social groups are more often suspected of AI use (kadoma2024generativeaiperceptualharms). Such consistent bias could disproportionately affect minorities and individuals seeking legitimate participation (liang23), such as non-native speakers, newcomers, and gender or racial minorities. It is concerning that most of our participants were confident in the efficacy of their heuristics and unaware of any potential bias. Our participants’ reliance on these heuristics suggests that moderators should be educated about potential bias and further studies are needed to evaluate the efficacy of detection heuristics in content moderation. Platforms may be able to help moderators by adding tools for AIGC detection. Earlier work demonstrated that “synthesized social signals,” or “social signals computationally derived from an account’s history,” help community members evaluate the credibility of other accounts (Im20). Similar tools may be useful for detecting AIGC. For example, showing information about a poster’s interactions with the system, such as time spent editing a post, could be helpful, though this also has user privacy implications. Surfacing a poster’s past behavior and other community memberships could also be relevant, as these have been shown to be useful in identifying misinformation posters (Bozarth2023; diakopoulos_diamonds_2010).
Given the combination of concerns about AIGC and the lack of tools to detect it, we expect to see communities develop their own, non-technical, means for demonstrating authenticity. For example, past research into anonymous online communities (bernstein14) found several such practices, such as triforcing (posting complex Unicode symbols only meaningful to the community) and timestamping (posting photos of oneself holding a piece of paper with the current date and time on it). These techniques signal in-group membership and authenticity by presenting something that only an authentic poster can produce. These forms of effortful communication (Zhang22) signal the poster’s effort, which makes them seem more valuable, given the relationship between effort and value mentioned by a participant in Section 4.1.1. In the age of low-effort AIGC we expect to see online communities place an even greater premium on effortful communication, which platform design can encourage. While these social innovations may emerge naturally from within a community, platforms can provide technical affordances to support their adoption, and the HCI research community can help.
@@ 6. Limitations and Future Work
Our qualitative study is based on interviews with moderators of communities that are concerned with AIGC use and are taking steps to curb it. This recruiting criteria supports our goal of understanding some of the attitudes and practical considerations of Reddit moderators who are adapting to AIGC. We note, however, that our study and results are not meant to provide a generalizable view of how all Reddit moderators and communities consider AIGC.
Our study purposefully omitted those communities that have fully embraced AIGC, as well as those that have not yet created rules governing its use. Our recruitment additionally focused on popular subreddits in order to target those that are currently most affected by AIGC. However, our participants often also moderated smaller subreddits and were able to share how their experiences and concerns are relevant to communities of various sizes. Additional studies could include moderators from a broader set of communities to provide a more complete picture of the range of ways that online communities are responding to AIGC, and to further explore the relationship between a community’s values (Prinster24) and its stance on AIGC. We hope that future quantitative work will complement our own, by investigating a wider range of communities and making stronger inferences about the platform as a whole. For example, future work could use surveys or computational content analysis to test how our qualitative findings generalize across subreddits of different size and type. Further, since AIGC is an internet-scale phenomenon, future studies should explore how moderators are responding on other platforms, which may have different purposes and norms.
@@ 7. Conclusion
We performed a qualitative investigation into volunteer Reddit moderators’ experiences with and attitudes towards AI-generated content. To better understand how CSCW researchers can support community self-determination we focused on communities that have restricted the use of AIGC. We heard from moderators that they are concerned about AIGC impacting the content quality, social dynamics, and governance processes in their communities, despite also seeing its potential to make it easier for some community members to contribute. In response, communities have enacted bans, which moderators enforce using time-intensive and imperfect heuristics. Despite enforcement challenges, our moderators report that, for now, their rules suffice, as they serve to clarify community norms about AIGC use.
As generative AI usage grows, the CSCW community needs to pay attention to how it affects the dynamics, practices, and health of online communities. There is much that the research community can do to develop tools and practices that may help community moderators guide their communities through this period of technological change. For example, new systems for effortful communication may help community members demonstrate authenticity, and new methods of communicating norms may help communities stay aligned. As a starting point, this paper documents the perspectives of moderators at this critical juncture, where, as one moderator commented, “there has to be a lot that we’re missing.”